I need your help on this problem:
4t^2+9=12t

Answer by stanbon(75887)

I need your help on this problem:
4t^2+9=12t
4t^2 - 12t + 9 = 0
---
Think of two numbers whose product is 4*9 = 36
and whose sum is -12.
--
The numbers are -6 and -6.
-----
Replace -12t by -6t-6t:
---
4t^2 - 6t -6t +9 = 0
---
Factor the 1st two and the last two terms separatel
2t(2t-3)-3(2t-3) = 0
Factor again:
(2t-3)(2t-3) = 0
t = 3/2 with multiplicity two.
==================================
Cheers,
Stan H.